For what values of k does 2x^2 - 5x - k = 0 have no real root.
Thanks so much in advance:)

Answer by TimothyLamb(4379)

---
2x^2 - 5x - k = 0
---
first, analyze this:
y(x) = 2x^2 - 5x = 0
---
the above quadratic equation is in standard form, with a=2, b=-5, and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 -5 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic vertex is a minimum at: ( x= 1.25, y= -3.125 )
---
now, analyze this:
2x^2 - 5x - k = 0
---
for (k < -3.125) the equation (2x^2 - 5x - k = 0) will have complex roots (i.e. non-real roots.)
NOTE:
-3.125 is the y-coordinate of the quadratic vertex above
---
check:
k < -3.125
arbitrarily set k = -3.1250000001
2x^2 - 5x + 3.1250000001 = 0
---
the above quadratic equation is in standard form, with a=2, b=-5, and c=3.1250000001
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 -5 3.1250000001
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two complex roots, i.e. the roots are no longer real for (k < -3.125):
x = 1.25 + (0.0000070710681)i
x = 1.25 - (0.0000070710681)i
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php